OpenStudy (anonymous):

WILL FAN AND MEDAL Estimate -square root 14 to the nearest integer

OpenStudy (anonymous):

3.741657386773941

OpenStudy (anonymous):

Used a calculator. :)

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

\[\sqrt{9}<\sqrt{14}<\sqrt{16}\] so it is near\[\sqrt{16}\] or 4

OpenStudy (anonymous):

yup!